please solve this problem
(17.92) to the power -1/9

Answer by lwsshak3(11628)

please solve this problem
(17.92) to the power -1/9
**
You can solve this problem directly on most scientific calculators like the TI-30.
or use logarithms.
Using logs:
log(17.92^(-1/9)=(-1/9)log(17.92)=(-1/9)*1.2533=-0.1393
convert to exponential form: base(10) raised to log of number(-0.1393)=number(17.92^(-1/9)
10^-.1393=.7256
..
From calculator:
17.92^(-1/9)=.7257